OpenStudy (anonymous):

what is the awnswer to the equation a/12=16

OpenStudy (unklerhaukus):

What do you get if you multiply both sides of the equation by twelve?

OpenStudy (anonymous):

144 and 192

OpenStudy (unklerhaukus):

Multiplication is the inverse operation operation of division, so the common factor will cancel on the left hand side of the equation \[\frac a{12}=16\\ 12\times\frac a{12}=12\times16\\ \color{red}{\cancel{\color{black}{12\times}}}\frac a{\color{red}{\cancel{\color{black}{12}}}}=12\times16\]